OpenStudy (anonymous):

You purchase 10,000 shares of a stock for $0.85 a share. One year later it is worth $0.95 a share. Find the percent of increase.

OpenStudy (anonymous):

(new cost - old cost) / old cost x 100

OpenStudy (anonymous):

\[\frac{ (new cost-oldcost) }{ oldcost }\times 100\]

OpenStudy (anonymous):

simply (.95-.85)/.85 *100=.1/.85 =11.76

OpenStudy (anonymous):

thank you very much